OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x cos x - 4 cos x sin2x -sin3x + 2 sin x cos x -sin2x + 2 sin x cos x 2 sin2x cos x - 2 sin x cos x

zepdrix (zepdrix):

Oh boy this one is a doozy :u

OpenStudy (anonymous):

haha does that mean its tough?

zepdrix (zepdrix):

Yah it's a bit of work to widdle it down :) \[\Large \cos(3x) \quad=\quad \cos(2x+x)\]Let's apply the `Angle Sum Identity for Cosine` \[\large \color{royalblue}{\cos(\alpha+\beta) \quad=\quad \cos \alpha \cos \beta - \sin \alpha \sin \beta}\]

OpenStudy (anonymous):

i honestly know didly squat about this one

zepdrix (zepdrix):

+_+

zepdrix (zepdrix):

Well look at the blue identity a sec. We'll apply it and our function will break down like this,\[\Large \cos(2x+x) \quad=\quad \cos(2x)\cos(x)-\sin(2x)\sin(x)\]Look at da blueee +_+ Understand how I applied that identity?

zepdrix (zepdrix):

That's not the answer by the way U: We have a long ways to go still.

OpenStudy (anonymous):

so we are just using identities?

zepdrix (zepdrix):

Yes

OpenStudy (anonymous):

ok yup i get that

zepdrix (zepdrix):

See how we have a cos(x) and a sin(x)? That's what we're trying to achieve. We want all of the angles to be in terms of just x. So we still need to deal with the cos(2x) and sin(2x). To do so, we can apply our `Double Angle Identities`:\[\large \color{ #CC0033}{\cos(2x) \quad=\quad 2\cos^2x-1}\]\[\large \color{#FF11AA}{\sin(2x)\quad=\quad 2\sin(x)\cos(x)}\]

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

\[\Large =\color{#CC0033}{\cos(2x)}\cos x-\color{#FF11AA}{\sin(2x)}\sin x\]So this becomes,\[\Large =\color{#CC0033}{\left(2\cos^2x-1\right)}\cos x-\color{#FF11AA}{\left(2\sin x \cos x\right)}\sin x\]

OpenStudy (anonymous):

ok cool got that

OpenStudy (anonymous):

then do i just facter those in?

zepdrix (zepdrix):

Yes, distribute the cosx and sinx respectively. We might have to fiddle with it a bit after that. There are `several` forms for the cosine double angle. So depending on which one THEY expected us to use, our answer might look a little drifferent. Multiply that all out, then we'll decide whether or not we need to use our Square Identity to fix it.

OpenStudy (anonymous):

ok one min im gunna east some dinner

OpenStudy (anonymous):

ok im back

OpenStudy (anonymous):

do you wanna see the answer choices that i have?

zepdrix (zepdrix):

No, I want you to do some of the work :3 lol Multiply that last step out.

OpenStudy (anonymous):

ok so i can replace the trig functions with a variable like u correct? while i do the algerbra

zepdrix (zepdrix):

\[\Large =\color{#CC0033}{\left(2\cos^2x-1\right)}\cos x-\color{#FF11AA}{\left(2\sin x \cos x\right)}\sin x\]Oh do something like this?\[\Large =\color{#CC0033}{\left(2u^2-1\right)}u-\color{#FF11AA}{\left(2vu\right)}v\]Yah I guess you could do that +_+

OpenStudy (anonymous):

ok so then i just take the 2u^2-1 and foil it to u correct? for the first step?

zepdrix (zepdrix):

mmmmmm ya :u

OpenStudy (anonymous):

ok so is it 2u^3-1u-2v^2u